# Equivariant Masked position prediction for efficient molecular representation

### v1

@@ 3.2.2 Position Prediction
It is important to note that operations within the channel dimension preserve overall equivariance. As a result, equation 14 satisfies the condition outlined in equation 9. Multi-channel features are utilized to retain high-frequency components. Prior work, as referenced in (Daigavane et al., 2024), attempts to normalize multi-channel features using the aggregation $\log(\sum_{channel}\exp(\cdot))$. Although this approach uses fewer parameters, the resulting distribution depends on the spherical harmonics representation, $\mathbf{f}_{k}$, which becomes sensitive after gridding. Ablation experiments in Appendix C.2 show that the distribution generated by equation 14 is more stable.
@@ 3.2.3 Loss Function
Radius loss. During training, the model is optimized by the distribution of true radius $||\vec{\mathbf{r}}_{ik}||$ and direction $\vec{\mathbf{r}}_{ik}/||\vec{\mathbf{r}}_{ik}||$, where $\vec{\mathbf{r}}_{ik}=\mathbf{p}_{i}-\mathbf{p}_{k}$. Intuitively, the deterministic vector $\vec{\mathbf{r}}_{ik}$ should correspond to a Dirac delta distribution denoted as $\delta(\vec{\mathbf{r}}-\vec{\mathbf{r}}_{ik})$. In practice, however, the model can learn the accurate $\vec{\mathbf{r}}_{ik}$ as long as the defined distribution can uniquely represent $\vec{\mathbf{r}}_{ik}$. The Dirac delta distribution can reduce training stability and lacks the ability to transfer to other conformations. To address this, we use a Gaussian distribution444 A common method to embedding distance Thomas et al. (2018); Zitnick et al. (2022) as a surrogate:
@@ 4 Related Work
Graph self-supervised methods. Leveraging the inherent graph structure of molecules, many graph-based self-supervised methods have the potential to train molecular models that capture transferable knowledge. For example, Hu et al. (2020) proposed graph context prediction and attribute masking methods to enhance molecular property prediction. GraphMAE (Hou et al., 2022) pre-trained molecular models using a generative decoder to reconstruct atomic and bond attributes. D-SLA (Kim et al., 2022) applied contrastive learning based on graph edit distance, improving predictions of molecular biochemical activities. Additionally, graph motifs—induced subgraphs that describe recurrence and significance—have increasingly been utilized to construct self-supervised learning frameworks for molecules (Rong et al., 2020; Zhang et al., 2021; Inae et al., 2023), facilitating the learning of multi-scale molecular information. However, despite their improvements in property prediction, these pre-training methods primarily focus on graph characteristics, such as geometric semantics and graph-level similarity, while neglecting the intrinsic quantum mechanisms within molecules. As a result, they are limited in their ability to predict molecular quantum mechanical properties. In this work, we propose EMPP, which introduces physical priors into learning tasks, enabling GNN models to capture key quantum mechanical features.
@@ 5.4 Ablation Study
To assess the impact of Gaussian mixture distributions in denoising, we applied the DP-TorchMD-Net Zaidi et al. (2023) to QM9 (HOMO, LUMO) as an auxiliary task, with the results displayed in Figure 3. By varying the standard deviation $\sigma$ of the Gaussian mixture, which controls the curvature, we observed changes in the model’s generalization performance. The experiment with $\sigma=0$ is equivalent to training without denoising. From Figure 3, we derive two key insights: (i) Without utilizing external data, the denoising method demonstrates limited potential to improve the model’s generalization ability, yielding results that are similar to or worse than the baseline without denoising. (ii) The performance of denoising is highly sensitive to the standard deviation, with different values leading to significantly divergent outcomes. In practical applications, because the curvature near local minima is often unknown, the optimal standard deviation typically requires empirical tuning, making it challenging to account for all possible local minima.

### v2

@@ 3.2.2 Position Prediction
It is important to note that operations within the channel dimension preserve overall equivariance. As a result, equation 14 satisfies the condition outlined in equation 9. Multi-channel features are utilized to retain high-frequency components. While EMPP’s method of masking atoms and predicting coordinates using spherical harmonic projection shares similarities with Symphony (Daigavane et al., 2024), their primary focuses diverge significantly: Symphony tackles molecular generation, whereas EMPP addresses a self-supervised task. This difference in objective leads to substantial variations in implementation, such as masking strategies and the design of the multi-channel spherical harmonic projection module. Detailed comparisons are provided in Appendix B.5.
@@ 3.2.3 Loss Function
Radius loss. During training, the model is optimized by the distribution of true radius $||\vec{\mathbf{r}}_{ik}||$ and direction $\vec{\mathbf{r}}_{ik}/||\vec{\mathbf{r}}_{ik}||$, where $\vec{\mathbf{r}}_{ik}=\mathbf{p}_{i}-\mathbf{p}_{k}$. Intuitively, the deterministic vector $\vec{\mathbf{r}}_{ik}$ should correspond to a Dirac delta distribution denoted as $\delta(\vec{\mathbf{r}}-\vec{\mathbf{r}}_{ik})$. In practice, however, the model can learn the accurate $\vec{\mathbf{r}}_{ik}$ as long as the defined distribution can uniquely represent $\vec{\mathbf{r}}_{ik}$. The Dirac delta distribution can reduce training stability and lacks the ability to transfer to other conformations. To address this, we use a Gaussian distribution444 A common method to embedding distance (Thomas et al., 2018; Zitnick et al., 2022) as a surrogate:
@@ 4 Related Work
Graph self-supervised methods. Leveraging the inherent graph structure of molecules, many graph-based self-supervised methods have the potential to train molecular models that capture transferable knowledge. For example, Hu et al. (2020) proposed graph context prediction and attribute masking methods to enhance molecular property prediction. GraphMAE (Hou et al., 2022) pre-trained molecular models using a generative decoder to reconstruct atomic and bond attributes. D-SLA (Kim et al., 2022) applied contrastive learning based on graph edit distance, improving predictions of molecular biochemical activities. Additionally, graph motifs—induced subgraphs that describe recurrence and significance—have increasingly been utilized to construct self-supervised learning frameworks for molecules (Rong et al., 2020; Zhang et al., 2021; Inae et al., 2023), facilitating the learning of multi-scale molecular information. These pre-training methods primarily focus on graph characteristics, while neglecting the intrinsic quantum mechanisms within molecules. As a result, they are limited in their ability to predict molecular quantum mechanical properties.
@@
Spherical harmonics projection. High-degree spherical harmonic representations with grid projections (Liao & Smidt, 2023; Passaro & Zitnick, 2023) have demonstrated significant capabilities in spatial description. Building on this, Symphony (Daigavane et al., 2024) introduced a neighbor-based spatial position prediction method, developing a framework for molecular generation. While EMPP shares some technical similarities with Symphony, their distinct tasks (molecular generation versus self-supervised learning) lead to differing implementation priorities. Symphony prioritizes flexibility in position prediction to ensure the sampling of diverse molecules, whereas EMPP emphasizes accuracy and well-posed position prediction to facilitate learning of genuine physical interactions.
@@ 5.4 Ablation Study
To assess the impact of Gaussian mixture distributions in denoising, we applied the DP-TorchMD-Net (Zaidi et al., 2023) to QM9 (HOMO, LUMO) as an auxiliary task, with the results displayed in Figure 3. By varying the standard deviation $\sigma$ of the Gaussian mixture, which controls the curvature, we observed changes in the model’s generalization performance. The experiment with $\sigma=0$ is equivalent to training without denoising. From Figure 3, we derive two key insights: (i) Without utilizing external data, the denoising method demonstrates limited potential to improve the model’s generalization ability, yielding results that are similar to or worse than the baseline without denoising. (ii) The performance of denoising is highly sensitive to the standard deviation, with different values leading to significantly divergent outcomes. In practical applications, because the curvature near local minima is often unknown, the optimal standard deviation typically requires empirical tuning, making it challenging to account for all possible local minima.
@@ B.4 Prediction with Neighbors
B.5 Differences from Symphony
Symphony (Daigavane et al., 2024) proposes a molecular generation framework based on equivariant representations. It constructs a fragment sequence by iteratively masking atoms, with each subsequent fragment masking one additional atom. Symphony then uses spherical harmonic projections of high-degree equivariant representations to encode the spatial distribution of atomic positions, enabling it to sample new atoms based on the predicted distribution. EMPP uses the similar technique to define position prediction module where the relative positions are projected onto a spatial distribution based on high-degree equivariant representations. However, EMPP and Symphony differ in several key aspects:
  • Masking Strategy: Symphony, geared towards molecular generation, masks all information about an atom, relying on the model’s prediction of a ”focus” atom to sample subsequent atoms. EMPP, conversely, masks only the position of a single atom, retaining the ground truth of its atomic type and neighbor information. EMPP’s position inference leverages all available real information from the molecule. These distinct strategies reflect their respective goals: Symphony prioritizes generating diverse molecules, while EMPP focuses on learning precise interaction patterns.
  • Prediction Horizon: Symphony predicts the position of an atom without considering the positions of subsequent atoms. EMPP, however, uses all available positional information except for the masked atom’s position. This difference again stems from their objectives: Symphony emphasizes diversity and fragment-based generation, while EMPP prioritizes accuracy by minimizing interfering factors in position prediction to capture genuine physical interactions.
  • Spatial Projection Module: Symphony normalizes multi-channel features using the aggregation $\log(\sum_{channel}\exp(\cdot))$. While parameter-efficient, this approach may make the resulting distribution sensitive to the spherical harmonic representation, $\mathbf{f}_{k}$, particularly after gridding. EMPP employs a neural network (detailed in equation 14) to aggregate multi-channel features on the sphere. Ablation experiments (Appendix C.2) demonstrate the greater stability of the distribution generated by EMPP’s neural network.
  • Label Definition: Symphony uses a Dirac delta distribution as the label for spatial projection to encourage sampling of physically plausible positions. EMPP, not designed for sampling, relaxes this constraint and adopts a smoother distribution to improve generalization performance.